# Enhancing inflationary model predictions via refined slow-roll dynamics

### v3

@@ Abstract
The inflationary paradigm not only addresses early universe puzzles but also aligns well with the observational constraints, with slow-roll inflationary models fitting the best. Evaluating these model predictions requires considering both slow-roll inflationary dynamics and the subsequent reheating epoch. This involves the quantitative analysis that takes into account the effective equation of state (EoS) and duration of reheating, connecting these with the perturbations generated deep during the inflationary era. Given the complexities involved, many approximations are often used for simplification. However, as future observations are expected to improve the accuracy of these observables significantly, this work takes a different approach. Instead of relying on approximations, and instead of looking into the complex effects of (pre-)reheating, we focus on the corrections arising purely from more accurate analytical evaluations of the perturbations generated during the inflationary era itself. This is because the reheating dynamics is model-dependent and lack a single concrete analytical description, and thus introduce large uncertainties, making robust predictions difficult. In this article, we mainly incorporate two improvements: the first is the accurate dynamics of the slow-roll evolution, and, thus, the end of inflation; and the second is the higher-order slow-roll corrections to the perturbed observables. Our findings indicate that, by implementing these corrections, the theoretical predictions improve significantly. It also indicates that seemingly minor corrections can have significant effects on the perturbed observables and these refined predictions can be compared with future observations to potentially rule out models and help resolve the degeneracy problem of the inflationary paradigm.
@@ I Introduction
The inflationary paradigm [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], which describes a brief period of accelerated expansion, is the most successful framework for explaining the dynamics of the early universe. Within this paradigm, slow-roll inflation involves field(s) slowly rolling down the potential, resulting in near-exponential expansion and providing an explanation for the observational constraints [32, 33, 34, 35]. As the field(s) approaches the bottom of the potential, inflation ends, and shortly afterward, the field(s) begins to oscillate around this minimum. This phase, known as the reheating epoch, involves the inflaton field(s) coupling with and decaying into other standard particles, leading to the onset of the radiation-dominated era [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58].
Theoretical constraints derived from an inflationary model, however, not only depend on the dynamics of inflation itself but also depend on the reheating era, and therefore, studying reheating dynamics is crucial for accurately predicting the early universe behavior. There are two main approaches to this: the first is the qualitative analysis [42, 36, 37, 38, 39, 41, 40, 43, 46] which examines the exact decay process of the inflaton field through the parametric resonance, leading to the understanding of the microphysics involved during the reheating epoch. The second approach, which is the main focus of this manuscript, is the quantitative analysis [59, 54, 60, 44, 61, 45, 47, 49, 62, 52, 63, 64]. This approach looks at the macrophysics of the reheating epoch by constraining the duration ($N_{\rm re}$) and the equation of state (EoS) parameter ($w_{\rm re}$) during the reheating era using perturbations generated deep in the early (inflation) universe.
In this analysis, the Hubble horizon and the associated length scales are used to constrain the duration when the pivot scale exits the Hubble horizon from the end of inflation, i.e., $N_{k}$, and the reheating parameters ($N_{\rm re},\ w_{\rm re}$). Given any model, using $N_{k}$, one can determine both the scalar and tensor perturbations associated with the pivot scale and compare this with the observations. Thus, the theoretically obtained perturbed observables, namely the scalar-spectral index $n_{s}$ and the tensor-to-scalar ratio $r$ are heavily dependent on the reheating parameters $N_{\rm re}$ and $w_{\rm re}.$ Since the perturbations ($n_{s}$ and $r$) are highly sensitive to $N_{k}$, and consequently, to ($N_{\rm re},\ w_{\rm re}),$ even small corrections or deviations in $N_{k}$ or reheating parameters can significantly improve $n_{s}.$ Ongoing experiments [32, 33, 34, 35] already measure $n_{s}$ with an accuracy of $\Delta n_{s}\sim 10^{-2}$ and upcoming experiments like PRISM [65], EUCLID [66], cosmic 21-cm surveys [67], and CORE experiments [68] are projected to improve this accuracy to $\Delta n_{s}\sim 10^{-3}$. Therefore, taking advantage of this, in this work, we focus on refining the analysis itself to obtain tighter constraints on the inflationary paradigm, as the analysis relies on numerous approximations. This refined approach, in turn, can help resolve the degeneracy problem of the inflationary paradigm [69, 25, 70, 71] by eliminating more inflation models.
Improving the analysis can again be performed in two main ways. The first one involves a detailed and careful consideration of the reheating analysis that can accurately depict the reheating parameters ($N_{\rm re},w_{\rm re}$). However, since it entails dealing with various distinct microphysical processes — requiring largely model-dependent treatment known as the qualitative analysis — and consequently introduces substantial uncertainties. For this reason, we have opted not to pursue this approach as it is beyond the scope of this work. Instead, we focus on the second, more straightforward improvement: the slow-roll methodology and approximations with a brief and simplest quantitative reheating epoch with constant $w_{\rm re}$. Recent findings in Refs. [72, 73] indicate that a near-accurate analytical solution can significantly enhance background predictions near the end of inflation, thereby improving $N_{k}$ [72]. For example, in the case of Starobinsky inflation, the improvement $\Delta N_{k}\sim 1.5$ suggests that
$n_{s}\sim 1-\frac{2}{N_{k}},\quad\Delta n_{s}\sim\frac{\Delta N_{k}}{N_{k}^{2}}\sim 10^{-3}.$
In addition to leading-order slow-roll approximations, higher-order slow-roll corrections [74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88] can be considered to further improve the evaluation of perturbed observables such as $n_{s}$ and $r$. Therefore, the objective of this work is to enhance the model predictions by implementing both the improved near-accurate dynamics and higher-order slow-roll corrections and comparing our results with the existing literature. As an example, we consider the most popular inflationary model: the Starobinsky model of inflation [1, 2, 8, 89, 90]. Using the conventional leading order analytical approximation, the constraint on $n_{s}$ is: $0.9613\leq n_{s}\leq 0.9641,$ which leads to further constrain $N_{\rm re}\leq 16$ (equivalently, the energy at the end of reheating, i.e., $T_{\rm re}\geq 10^{10}$ GeV, assuming $w_{\rm re}=0$). However, after implementing the accurate dynamics as well as the higher-order slow-roll corrections, the bounds on these parameters change significantly, and the new constraint on $n_{s}$ for numerical solution becomes: $0.9613\leq n_{s}\leq 0.9649,$ which leads to further constrain $N_{\rm re}\leq 21$ (and subsequently, $T_{\rm re}\geq 10^{8}$ GeV), while for higher-order corrections these become: $0.9613\leq n_{s}\leq 0.9653$, constraints on reheating parameters are, $N_{\rm re}\leq 24$ and $T_{\rm re}\geq 10^{7}$, which has the potential to rule out the model from future observations. The result is unexpected, as one might anticipate the corrections to be negligible. Thus, instead of relying on the conventional approximation method, our results demonstrate the critical need for corrections due to the accurate dynamics in each model of inflation, which may lead to a significant change in theoretical predictions and potentially rule out models from observations. This implies that even seemingly minor improvements to the background dynamics can have a substantial impact on the theoretically obtained perturbed observables, which must not be ruled out without verification.
The article is structured as follows: the following section II presents the model and the subsequent dynamics for the early universe dynamics. In this section, we also demonstrate the slow-roll conditions as well as the slow-roll approximations. In the subsequent section III, we study the quantitative analysis of reheating linking the inflationary perturbed variables to the observable constraints. This section, as an example, also examines the popular Starobinsky model of inflation, and with the help of the reheating analysis along with leading-order slow-roll approximations, we carefully obtain the observational consequences of the model. In Sec. IV, we highlight the dire need for the corrections to the dynamics, subsequently implement these changes, and obtain the improved constraint of the model, which turns out to be significant. Finally, we conclude our work in Sec. V.
@@
A brief explanation about some terms and symbols we’ll be using. In our work, we use natural units where $\hbar=c=k_{B}=1$. We also define the reduced Planck mass as $M_{\rm pl}\equiv(8\pi G)^{-1/2}=1$. The metric signature we use is (,+,+,+)(-,+,+,+). Greek indices are contracted with the metric tensor $g_{\mu\nu}$, while Latin indices will be contracted with the Kronecker delta $\delta_{ij}$. We use symbols $\partial$ and $\nabla$ for partial and covariant derivatives, respectively. Additionally, the dots and primes above symbols represent derivatives for cosmic and conformal time, respectively, related to the Friedmann-Lemaître-Robertson-Walker (FLRW) line-element.
@@ II General equations and slow-roll conditions
In this section, we discuss the general equations and solutions for a slow-roll inflationary universe. For that, let us consider the simplest action with a single canonical scalar field $\phi$ minimally coupled to gravity with a potential $V(\phi)$, which can be written as
@@
$H=\sqrt{\frac{V}{3-\frac{1}{2}\phi_{N}^{2}}},\qquad\frac{H_{N}}{H}=-\frac{1}{2}\phi_{N}^{2},$ (9)
@@
where, $A_{xx}=\partial^{2}A/\partial{x}^{2}.$ Given a potential $V(\phi)$, one can solve the above equations to correctly obtain the dynamics of the early universe. Once the background equations are solved, one can in principle, solve the scalar $(\mathcal{R}_{k})$ and tensor $(h_{k})$ perturbations and evaluate the power spectra:
@@
are the scalar and tensor spectral index and and
@@
is defined as the tensor-to-scalar ratio, which we compare with the observations. The current observations (PLANCK [32, 33, 35], BICEP/Keck [34] ) suggest that, at the pivot scale ($k_{*}=0.05\,\text{Mpc}{}^{-1}$), the amplitude of the scalar power spectrum is $A_{\mathcal{R}}\simeq 2.101^{+0.031}_{-0.034}\times 10^{-9}(68\%\,\,\text{CL})$ with the scalar spectral index being $n_{{}_{\rm s}}=0.9672\pm 0.0059\,(68\%\,\,\text{CL})$, while the ratio $(r)$ of the tensor-to-scalar perturbation amplitudes is bounded by $r<0.028\,(95\%\,\,\text{CL})$. These observations agree remarkably well with the predictions of a near-scale-invariant spectrum of density perturbations in the inflationary paradigm. In the next section, we will discuss in detail, how, without explicitly evaluating the perturbations, one can obtain these observables using the background solutions in the case of slow-roll inflation.
@@
Let us first define the two slow-roll parameters $\epsilon_{1}$ and $\epsilon_{2}$ as
@@
The first variable $\epsilon_{1}<1$ ensures that the universe is accelerating. On the other hand, the second variable $\epsilon_{2}\ll 1$ ensures the sufficient duration of the inflationary phase. In the case of slow-roll, both these parameters are extremely small, i.e.,
@@
These two condition are also referred to as slow-roll conditions. Under this conditions, Eqs. (9) and (10) become
@@
and the slow-roll parameters can be re-expressed in terms of shape of potential as
@@
The above expressions are obtained using leading-order slow-roll approximation, assuming the higher-order contributions are heavily suppressed. Please note that these observables depend on the time variable $N_{k}$, representing the duration between the end of inflation and the epoch when the pivot scale $k$ leaves the Hubble horizon. Typically, this duration is considered to be $N_{k}\sim 50-60$. However, it is a model-dependent quantity and in order to obtain the exact duration $N_{k}$, one must consider the post-inflationary dynamics, i.e., the reheating epoch. In the very next section, we will focus on this, which is also known as the quantitative analysis of reheating.
@@ III Quantitative analysis of reheating
This epoch involves intricate microphysical processes and is largely model-dependent. As a result, any analytical extension of its dynamics carries substantial uncertainties, making it extremely challenging to describe the epoch through a single, unified analytical framework. To circumvent this difficulty and achieve a model-independent treatment, a simplified quantitative approach has been developed in which the equation-of-state parameter during the entire reheating epoch, i.e., $w_{\rm re}$ is assumed to remain constant. Although this approximation does not capture the full microphysical complexity of reheating, it provides a useful macrophysical description that successfully connects inflationary predictions with observations and thereby enables broad, indirect constraints on the reheating scenario. In this case, the energy density during reheating effectively behaves as
@@
We also define the duration of reheating $N_{\rm re}$ as
@@
where $a_{\rm end},\ a_{\rm re}$ are the scale factor solutions at the end of the inflation and reheating epoch, respectively. In order to calculate the duration $N_{\rm re}$, let us compare the mode of interest with the present scale as
@@
where $N_{k}\equiv{\rm ln}\left(\frac{a_{k}}{a_{\rm end}}\right),\,N_{\rm RD}\equiv{\rm ln}\left(\frac{a_{\rm re}}{a_{\rm eq}}\right).$ At the end of reheating, the energy density is
@@
$N_{\rm re}=\frac{4}{1-3w_{\rm re}}\left(-N_{k}-\frac{1}{4}\ln\left(\frac{30}{\pi^{2}g_{\rm{re}}}\right)-\frac{1}{3}\ln\left(\frac{11g_{\rm{s,re}}}{43}\right)-\ln\left(\frac{k}{a_{0}T_{0}}\right)-\frac{1}{4}\ln\rho_{\rm{end}}+\ln H_{k}\right).$
@@
$H_{k}=\frac{\pi\sqrt{rA_{\mathcal{R}}}}{\sqrt{2}}.$ (26)

### v4

# Precision Inflationary Predictions: Impact of Accurate End-of-Inflation Dynamics

@@ Abstract
The precision era of cosmology demands accurate theoretical predictions from inflationary models. In quantitative reheating analyses, inflationary observables depend sensitively on the number of e-folds between horizon exit and the end of inflation, $N_{k}$, whose determination relies on slow-roll approximations near the end of inflation. Since inflation ends when the first slow-roll parameter reaches unity, even modest inaccuracies in this approximation can shift the end of inflation and thereby alter $N_{k}$, leading to modifications in predicted observables — including those evaluated at leading order. While such effects are implicit in standard treatments, their quantitative impact on observable constraints has not been systematically assessed. In this work, we first re-evaluate leading-order slow-roll predictions using an improved determination of $N_{k}$ within a simple quantitative reheating framework, and then incorporate higher-order slow-roll corrections consistently with the revised background evolution. Applying this framework to the Starobinsky model, we find that improved end-of-inflation dynamics alone can induce shifts of order $\Delta n_{s}\sim 10^{-3}$, while higher-order slow-roll corrections provide additional refinements at the $\sim 4\times 10^{-4}$ level. The cumulative effect results in a maximum shift of $\Delta n_{s}\sim 1.2\times 10^{-3}$ within the allowed reheating range. These results demonstrate that an accurate determination of the end of inflation is essential for precision tests of inflationary models.
@@ I Introduction
The inflationary paradigm [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] provides a compelling description of the early universe, resolving the horizon and flatness problems while generating primordial perturbations in excellent agreement with observations of the cosmic microwave background and large-scale structure. In particular, slow-roll inflationary models predict an almost scale-invariant spectrum of fluctuations consistent with current data [32, 33, 34, 35]. As cosmological observations steadily improve in precision, however, the reliability of inflationary predictions increasingly depends on the accuracy of the underlying theoretical framework. In this precision-driven context, effects that are traditionally regarded as subleading can become observationally relevant, motivating a reassessment of standard approximation schemes.
Connecting inflationary predictions to observations requires accounting for the post-inflationary evolution of the universe, in particular the reheating epoch that bridges inflation and the radiation-dominated era [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. In quantitative analyses [59, 54, 60, 44, 61, 45, 47, 49, 62, 52, 63, 64], reheating is commonly characterized by an effective equation-of-state parameter and the duration of the reheating phase, which together determine the number of e-folds between horizon exit of observable modes and the end of inflation. Since inflationary observables, such as the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$, are highly sensitive to this duration, uncertainties in reheating directly propagate into predicted cosmological parameters.
Despite the sensitivity of inflationary predictions to post-inflationary evolution, it is commonly assumed that leading-order slow-roll approximations provide sufficiently accurate estimates of the inflationary background dynamics, including the determination of the end of inflation. In standard treatments, the end of inflation is inferred from slow-roll conditions, although inflation strictly terminates when the first slow-roll parameter reaches unity. Even modest inaccuracies in this approximation shift the end of inflation and therefore modify the duration of inflation. Since the number of e-folds $N_{k}$ between horizon exit and the end of inflation depends directly on this endpoint, a careful and consistent determination of the end of inflation becomes essential for reliable theoretical predictions.
The central role of $N_{k}$ becomes particularly evident in quantitative reheating analyses, where it appears explicitly in the consistency relation that connects the comoving pivot scale to the post-inflationary expansion history. More precisely, $N_{k}$ depends on both the effective reheating equation-of-state parameter $w_{\rm re}$ and the reheating duration $N_{\rm re}$, subject to the physical constraint $N_{\rm re}\geq 0$, where $N_{\rm re}\ll 1$ corresponds to instantaneous reheating. Through this relation, $N_{k}$ determines the value of the inflaton field at horizon exit and hence the slow-roll parameters that enter predictions for observables such as the scalar spectral index $n_{s}$ and the tensor-to-scalar ratio $r$. Even within leading-order slow-roll expressions, small shifts in $N_{k}$ directly translate into corrections to $n_{s}$ and $r$ at a level comparable to current [32, 33, 34, 35] and forthcoming observations such as PRISM [65], EUCLID [66], cosmic 21-cm surveys [67], and CORE [68] sensitivities. As cosmological measurements continue to improve in precision, achieving reliable theoretical control over the determination of $N_{k}$ becomes increasingly important. However, in most standard treatments, the effect of accurately determining the end of inflation itself is implicitly assumed to be negligible and has not been systematically quantified.
Motivated by this sensitivity, the primary objective of this work is to obtain a more reliable determination of the duration of inflation and the corresponding number of e-folds $N_{k}$ by improving the treatment of the inflationary background dynamics near the end of inflation. Since inflation terminates when the first slow-roll parameter reaches unity, even modest inaccuracies in the slow-roll approximation can shift the end of inflation and thereby alter $N_{k}$, leading to modifications in inflationary observables — including those evaluated at leading order. While recent studies [72, 73] have indicated that improved treatments of the background evolution can induce non-negligible corrections to the estimation of $N_{k}$, these effects have not been systematically quantified with explicit numerical implications for inflationary observables. In this work, we address this gap by carrying out a detailed analysis in two stages. First, we reassess leading-order slow-roll predictions using an improved determination of $N_{k}$ within a simple and controlled quantitative reheating framework, thereby isolating the impact of end-of-inflation corrections on baseline theoretical predictions. Second, we incorporate higher-order slow-roll corrections [74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88] consistently with the revised background evolution to evaluate additional refinements. As a representative and observationally viable benchmark, we apply this framework to the Starobinsky model of inflation [1, 2, 8, 89, 90] and demonstrate that corrections often regarded as negligible can lead to appreciable modifications of inflationary predictions in the precision cosmology era.
This article is organized as follows. In Sec. II, we introduce the inflationary model and the corresponding background dynamics, including the slow-roll conditions and the associated approximations. Sec. III presents the quantitative reheating analysis that connects inflationary perturbations to observable constraints, where the Starobinsky model is examined as a representative example using leading-order slow-roll expressions. In Sec. IV, we implement corrections to the inflationary dynamics and assess their impact on theoretical predictions and reheating constraints. Our conclusions are summarized in Sec. V.
@@
Throughout this work, we employ natural units with $\hbar=c=k_{B}=1$ and set the reduced Planck mass to $M_{\rm pl}\equiv(8\pi G)^{-1/2}=1$. The metric signature is (,+,+,+)(-,+,+,+), Greek indices are contracted with the spacetime metric $g_{\mu\nu}$, and Latin indices with the Kronecker delta $\delta_{ij}$. We denote partial and covariant derivatives by $\partial$ and $\nabla$, respectively, while overdots and primes represent derivatives with respect to cosmic and conformal time in the Friedmann–Lemaître–Robertson–Walker background.
@@ II General equations and slow-roll conditions
In this section, we present the general background equations governing single-field inflation and review the standard slow-roll framework that underlies conventional analytical predictions. This formulation will serve as the baseline against which we later assess corrections arising from a more accurate treatment of the end of inflation. For that, let us consider the simplest action with a single canonical scalar field $\phi$ minimally coupled to gravity with a potential $V(\phi)$, which can be written as
@@
$H^{2}=\frac{V}{3-\frac{1}{2}\phi_{N}^{2}},\qquad\frac{H_{N}}{H}=-\frac{1}{2}\phi_{N}^{2},$ (9)
@@
where, $A_{xx}=\partial^{2}A/\partial{x}^{2}.$ For a given potential $V(\phi)$, these equations fully determine the background dynamics of inflation. In conventional analyses, these equations are typically simplified under the slow-roll approximation, yielding analytical expressions for inflationary observables. In the following, we briefly review this standard framework, which will serve as the baseline for the improved treatment developed later in this work.
Once the background equations are solved, one can in principle, solve the scalar $(\mathcal{R}_{k})$ and tensor $(h_{k})$ perturbations and evaluate the power spectra:
@@
are the scalar and tensor spectral indices and
@@
is defined as the tensor-to-scalar ratio. The current observations constrain $A_{\mathcal{R}}\simeq 2.101^{+0.031}_{-0.034}\times 10^{-9}(68\%\,\,\text{CL})$, $n_{{}_{\rm s}}=0.9672\pm 0.0059\,(68\%\,\,\text{CL})$ and $r<0.028\,(95\%\,\,\text{CL})$ at $k_{*}=0.05\,\text{Mpc}{}^{-1}$ (PLANCK [32, 33, 35], BICEP/Keck [34] ). These observations agree remarkably well with the predictions of a near-scale-invariant spectrum of density perturbations in the inflationary paradigm. In the next section, we will discuss in detail, how, without explicitly evaluating the perturbations, one can obtain these observables using the background solutions in the case of slow-roll inflation.
@@
We adopt the Hubble-flow definition of slow-roll parameters. Let us first define the two slow-roll parameters $\epsilon_{1}$ and $\epsilon_{2}$ as
@@
While $\epsilon_{1}<1$ ensures that the universe is accelerating, the condition $\epsilon_{2}\ll 1$ ensures the sufficient duration of the inflationary phase. Inflation ends when the first slow-roll parameter reaches unity, i.e., $\epsilon_{1}=1$, without invoking slow-roll approximations. In the case of slow-roll, both these parameters are extremely small, i.e.,
@@
These two conditions are also referred to as the slow-roll conditions. Under this conditions, Eqs. (9) and (10) become
@@
and, using the above equation, the slow-roll parameters can be re-expressed in terms of shape of potential as
@@
The above expressions are obtained using leading-order slow-roll approximation, assuming the higher-order contributions are heavily suppressed. These observables depend explicitly on the e-folding number $N_{k}$, defined as the interval between horizon exit of the pivot scale and the end of inflation. Since $N_{k}$ is determined by the precise end of inflation, any inaccuracy in the slow-roll estimation of the end-of-inflation propagates directly into these leading-order predictions. Typically, this duration is considered to be $N_{k}\sim 50-60$. However, as we shall demonstrate, even shifts of order unity in $N_{k}$ can translate into corrections in $n_{s}$ at the level of $10^{-3}$, which is comparable to present and projected observational precision. However, it is a model-dependent quantity and in order to obtain the exact duration $N_{k}$, one must consider the post-inflationary dynamics, i.e., the reheating epoch. A consistent determination of $N_{k}$ therefore requires incorporating the post-inflationary expansion history. In the next section, we implement this through the quantitative analysis of reheating.
@@ III Quantitative analysis of reheating
This epoch involves intricate microphysical processes and is largely model-dependent. As a result, any analytical extension of its dynamics carries substantial uncertainties, making it extremely challenging to describe the epoch through a single, unified analytical framework. To circumvent this difficulty and achieve a model-independent treatment, a simplified quantitative approach has been developed in which the equation-of-state parameter during the entire reheating epoch, i.e., $w_{\rm re}$ is assumed to remain constant. Although this approximation does not capture the full microphysical complexity of reheating, it provides a useful average macrophysical description that successfully connects inflationary predictions with observations and thereby enables broad, indirect constraints on the reheating scenario. In this case, the energy density during reheating effectively behaves as
@@
This relation follows from energy conservation $\dot{\rho}+3H(1+w)\rho=0$. We also define the duration of reheating $N_{\rm re}$ as
@@
where $a_{\rm end},\ a_{\rm re}$ are the scale factor solutions at the end of the inflation and reheating epoch, respectively. Physical consistency requires $N_{\rm re}\geq 0,$ where $N_{\rm re}\ll 1$ corresponds to instantaneous reheating. In order to calculate the duration $N_{\rm re}$, let us compare the mode of interest with the present scale as
@@
where $N_{k}\equiv{\rm ln}\left(\frac{a_{\rm end}}{a_{k}}\right),\,N_{\rm RD}\equiv{\rm ln}\left(\frac{a_{\rm eq}}{a_{\rm re}}\right).$ At the end of reheating, the energy density is
@@
$N_{\rm re}=\frac{4}{1-3w_{\rm re}}\left(-N_{k}-\frac{1}{4}\ln\left(\frac{30}{\pi^{2}g_{\rm{re}}}\right)-\frac{1}{3}\ln\left(\frac{11g_{\rm{s,re}}}{43}\right)-\ln\left(\frac{k}{a_{0}T_{0}}\right)-\frac{1}{4}\ln\left(\frac{\rho_{\rm{end}}}{H_{k}^{4}}\right)\right).$
@@
$H_{k}=\pi\sqrt{\frac{rA_{\mathcal{R}}}{2}}.$ (26)